# Prospects for detecting charged long-lived BSM particles at MoEDAL-MAPP experiment: A mini-review

### v1

@@
Abstract
@@ 1 Introduction
The search for charged LLPs has become a key frontier at the LHC. The general-purpose detectors, ATLAS and CMS, have deployed extensive search programs targeting Heavy Stable Charged Particles (HSCPs) and multi-charged objects. These analyses typically exploit anomalous ionisation energy loss ($dE/dx$) in the inner trackers and Time-of-Flight (ToF) measurements in the muon systems to identify slow-moving, heavy candidates. Recently, the ATLAS collaboration published results using the full Run 2 dataset ($139\penalty 10000\ \text{fb}^{-1}$), placing stringent limit on meta-stable gluino R-hadrons, $m\gtrsim 2$ TeV [ATLAS:2022pib], as well as multi-charged particles ($|q|\geq 2e$) up to the TeV range [ATLAS:2023zxo]. Similarly, the CMS collaboration has published its updated HSCP search using the Run 2 data ($101\penalty 10000\ \text{fb}^{-1}$), excluding gluino R-hadrons with masses up to $\sim 2$ TeV and identifying exclusion limit for a doubly charged fermion $\gtrsim 1.4$ TeV [CMS:2024nhn].
@@
The MoEDAL experiment is able to cover some of these blind spots. MoEDAL employs passive Nuclear Track Detectors (NTDs) and Magnetic Monopole Trappers (MMTs) that are sensitive to highly ionizing particles without the need for electronic triggers or precise timing [MoEDAL:2009jwa]. This allows for the detection of particles with arbitrarily low velocities and very high electric or magnetic charges, a regime often inaccessible to ATLAS and CMS. Specifically, the MoEDAL collaboration has searched for Highly Electrically Charged Objects (HECOs) with $|q|\geq 10e$ using Run 1 [MoEDAL:2021mpi] and Run 2 [MoEDAL:2023ost] data.
@@ 2 MoEDAL-MAPP experiment
The other detector subsystem is MMT, designed to capture magnetic monopoles for later examination, and a pixel detector, TPX, for real-time monitoring of the radiative environment. In preparation for the Run 3 data taking period, MoEDAL has been augmented with a MoEDAL Apparatus for Penetrating Particles (MAPP), a scintillator detector for discovery of Feebly Interacting Particles. The collaboration decided on changing its name and has been using "MoEDAL-MAPP" since then.
@@ 3 Methodology of phenomenological studies
Figure 1: Nuclear Track Detector panel placement for the MoEDAL Run 2 geometry. Image taken from [Maselek:2023fvy].
@@ 4.1 Proof-of-concept: neutralinos decaying to meta-stable staus
Figure 2: Velocity distributions for 1 TeV staus, Higgsinos, and gluinos produced via Drell-Yan process. Plot taken from [Felea:2020cvf].
@@
Figure 3: Expected sensitivity of the MoEDAL experiment to supersymmetric model. Red and blue contours correspond to Run 2 and "ideal" MoEDAL geometries, respectively. Solid line marks contours for $\rm N_{sig}=1$, while dashed lines are for $\rm N_{sig}=2$. The yellow contour shows constraints from ATLAS search for HSCPs [ATLAS:2019gqq], done for $L=36.1\penalty 10000\ {\rm fb}^{-1}$. The orange contour is ATLAS constraint projected for the end of the Run 3 data taking period, $L=300\penalty 10000\ {\rm fb}^{-1}$. Plot taken from [Felea:2020cvf].
@@ 4.2 Searches for supersymmetric long-lived particles
Table 1: Expected MoEDAL’s sensitivity to long-lived supersymmetric particles at the end of Run 3 ($L=30\penalty 10000\ {\rm fb}^{-1}$) compared with ATLAS and CMS constraints (in parentheses). Constraints for squarks and Higgsinos (in double parentheses) were obtained by recasting results from Ref. [ATLAS:2019gqq]. Table taken from [Acharya:2020uwc].
@@
Table 2: Expected MoEDAL’s sensitivity to long-lived doubly-charged particles at the end of Run 3 ($L=30\penalty 10000\ {\rm fb}^{-1}$) compared with CMS constraints for singlet fermions (in parentheses) from Ref. [CMS:2016kce] and reinterpreted results (in double parentheses) for other types of particles. Table taken from [Acharya:2020uwc].
@@ 4.3 Testing radiative neutrino mass models
Table 3: Model-independent mass limits on long-lived multiply charged colour-singlet BSM particles. The first column contains limits for fermions from Ref. [ATLAS:2018imb] for $L=36\penalty 10000\ {\rm fb}^{-1}$, and our recast for scalar particles (in double parentheses). The second column contains the projection for fermions for Run 3 $L=300\penalty 10000\ {\rm fb}^{-1}$, taken from [Jager:2018ecz]. The last two columns contain the expected sensitivity of MoEDAL for $L=30\penalty 10000\ {\rm fb}^{-1}$ and $L=300\penalty 10000\ {\rm fb}^{-1}$, respectively, for $\rm N_{sig}=3$ ($\rm N_{sig}=1$). All numbers are in GeV. Table taken from [Hirsch:2021wge].
@@
Table 4: Model-independent mass limits on long-lived multiply charged colour-(anti)triplet scalar particles. The first column contains limits (in double parentheses) estimated in Ref. [Jager:2018ecz], by scaling up 7 and 8 TeV CMS results [CMS:2013czn] to 13 TeV and $L=30\penalty 10000\ {\rm fb}^{-1}$. The second column contains projections at the end of Run 3 ($L=300\penalty 10000\ {\rm fb}^{-1}$), taken from [Jager:2018ecz]. The last two columns contain the expected sensitivity of MoEDAL for $L=30\penalty 10000\ {\rm fb}^{-1}$ and $L=300\penalty 10000\ {\rm fb}^{-1}$, respectively, for $\rm N_{sig}=3$ ($\rm N_{sig}=1$). All numbers are in GeV. Table taken from [Hirsch:2021wge].
@@ 4.4 Searches for multiply charged BSM particles
Figure 4: Expected sensitivity of LHC searches for multiply charged long-lived particles at the end of HL-LHC. Results are for colour-singlet scalars (top left), colour-singlet fermions (top right), colour-triplet scalars (bottom left), and colour-triplet fermions (bottom right). Figure taken from [Altakach:2022hgn].

### v2

@@
[…]
Prospects for detecting charged long-lived BSM particles at MoEDAL-MAPP experiment:
A mini-review
[…]
Streszczenie
@@ 1 Introduction
The search for charged LLPs has become a key frontier at the LHC. The general-purpose detectors, ATLAS and CMS, have deployed extensive search programs targeting Heavy Stable Charged Particles (HSCPs) and multi-charged objects. These analyses typically exploit anomalous ionisation energy loss ($dE/dx$) in the inner trackers and Time-of-Flight (ToF) measurements in the muon systems to identify slow-moving, heavy candidates. Recently, the ATLAS collaboration published results using the full Run 2 dataset ($139\penalty 10000\ \text{fb}^{-1}$), placing stringent limit on meta-stable gluino $R$-hadrons, $m\gtrsim 2$ TeV [ATLAS:2022pib], as well as multi-charged particles with electric charges in the range $|q|\in[2e,7e]$, excluding masses up to $1.6$ TeV for $|q|=6e$ [ATLAS:2023zxo]. Similarly, the CMS collaboration has updated its HSCP search using Run 2 data ($101\penalty 10000\ \text{fb}^{-1}$), identifying exclusion limits for doubly charged fermions $\geq 1.4$ TeV [CMS:2024nhn].
@@
The MoEDAL experiment is able to cover some of these blind spots. MoEDAL employs passive Nuclear Track Detectors (NTDs) and Magnetic Monopole Trappers (MMTs) that are sensitive to highly ionizing particles without the need for electronic triggers or precise timing [MoEDAL:2009jwa]. This allows for the detection of particles with arbitrarily low velocities and very high electric or magnetic charges, a regime often inaccessible to ATLAS and CMS. Specifically, the MoEDAL collaboration has searched for Highly Electrically Charged Objects (HECOs) with $|q|\in[15e,175e]$ using Run 1 [MoEDAL:2021mpi] ($E_{\rm CM}=8\penalty 10000\ {\rm TeV},\penalty 10000\ L=2.2\penalty 10000\ {\rm pb}^{-1}$) data, and was able to put mass limits ranging from 70 to 1020 GeV, depending on the electric charge and spin (see Tab. 2 in Ref. [MoEDAL:2021mpi]). The most recent search was done on Run 2 [MoEDAL:2023ost] ($E_{\rm CM}=13\penalty 10000\ {\rm TeV},\penalty 10000\ L=6.46\penalty 10000\ {\rm fb}^{-1}$) data, where HECOs with charges $q\in[10e,400e]$ were considered, and the limits spanned from 80 GeV to $\approx$ 3.9 TeV (see Tab. II in Ref. [MoEDAL:2023ost]).
@@ 2 MoEDAL-MAPP experiment
The other detector subsystem is MMT, designed to capture magnetic monopoles for later examination, and a pixel detector, TPX, for real-time monitoring of the radiative environment. In preparation for the Run 3 data taking period, MoEDAL has been augmented with a MoEDAL Apparatus for Penetrating Particles (MAPP)[Kalliokoski:2023cgw], a scintillator detector for discovery of Feebly Interacting Particles. The collaboration decided on changing its name and has been using "MoEDAL-MAPP"since then.
@@ 3 Methodology of phenomenological studies
Rysunek 1: Nuclear Track Detector panel placement for the MoEDAL Run 2 geometry. Image taken from [Maselek:2023fvy].
@@ 4.1 Proof-of-concept: neutralinos decaying to meta-stable staus
Rysunek 2: Velocity distributions for 1 TeV staus, Higgsinos, and gluinos produced via Drell-Yan process. Plot taken from [Felea:2020cvf].
@@
Rysunek 3: Expected sensitivity of the MoEDAL experiment to supersymmetric model. Red and blue contours correspond to Run 2 and "ideal"MoEDAL geometries, respectively. Solid line marks contours for $\rm N_{sig}=1$, while dashed lines are for $\rm N_{sig}=2$. The yellow contour shows constraints from ATLAS search for HSCPs [ATLAS:2019gqq], done for $L=36.1\penalty 10000\ {\rm fb}^{-1}$. The orange contour is ATLAS constraint projected for the end of the Run 3 data taking period, $L=300\penalty 10000\ {\rm fb}^{-1}$. Plot taken from [Felea:2020cvf].
@@ 4.2 Searches for supersymmetric long-lived particles
Tabela 1: Expected MoEDAL’s sensitivity to long-lived supersymmetric particles at the end of Run 3 ($L=30\penalty 10000\ {\rm fb}^{-1}$) compared with ATLAS and CMS constraints (in parentheses). Constraints for squarks and Higgsinos (in double parentheses) were obtained by recasting results from Ref. [ATLAS:2019gqq]. Table taken from [Acharya:2020uwc].
@@
Tabela 2: Expected MoEDAL’s sensitivity to long-lived doubly-charged particles at the end of Run 3 ($L=30\penalty 10000\ {\rm fb}^{-1}$) compared with CMS constraints for singlet fermions (in parentheses) from Ref. [CMS:2016kce] and reinterpreted results (in double parentheses) for other types of particles. Table taken from [Acharya:2020uwc].
@@ 4.3 Testing radiative neutrino mass models
Tabela 3: Model-independent mass limits on long-lived multiply charged colour-singlet BSM particles. The first column contains limits for fermions from Ref. [ATLAS:2018imb] for $L=36\penalty 10000\ {\rm fb}^{-1}$, and our recast for scalar particles (in double parentheses). The second column contains the projection for fermions for Run 3 $L=300\penalty 10000\ {\rm fb}^{-1}$, taken from [Jager:2018ecz]. The last two columns contain the expected sensitivity of MoEDAL for $L=30\penalty 10000\ {\rm fb}^{-1}$ and $L=300\penalty 10000\ {\rm fb}^{-1}$, respectively, for $\rm N_{sig}=3$ ($\rm N_{sig}=1$). All numbers are in GeV. Table taken from [Hirsch:2021wge].
@@
Tabela 4: Model-independent mass limits on long-lived multiply charged colour-(anti)triplet scalar particles. The first column contains limits (in double parentheses) estimated in Ref. [Jager:2018ecz], by scaling up 7 and 8 TeV CMS results [CMS:2013czn] to 13 TeV and $L=30\penalty 10000\ {\rm fb}^{-1}$. The second column contains projections at the end of Run 3 ($L=300\penalty 10000\ {\rm fb}^{-1}$), taken from [Jager:2018ecz]. The last two columns contain the expected sensitivity of MoEDAL for $L=30\penalty 10000\ {\rm fb}^{-1}$ and $L=300\penalty 10000\ {\rm fb}^{-1}$, respectively, for $\rm N_{sig}=3$ ($\rm N_{sig}=1$). All numbers are in GeV. Table taken from [Hirsch:2021wge].
@@ 4.4 Searches for multiply charged BSM particles
Rysunek 4: Expected sensitivity of LHC searches for multiply charged long-lived particles at the end of HL-LHC. Results are for colour-singlet scalars (top left), colour-singlet fermions (top right), colour-triplet scalars (bottom left), and colour-triplet fermions (bottom right). Figure taken from [Altakach:2022hgn].